OpenStudy (anonymous):

A quadratic equation is shown below: 9x2 - 16x + 60 = 0 Part A: Describe the solution(s) to the equation by just determining the discriminant. Show your work. (3 points) Part B: Solve 4x2 + 8x - 5 = 0 by using an appropriate method. Show the steps of your work and explain why you chose the method used. (4 points) Will medal fan and write a testimony.. please help! Part C: Solve 2x2 -12x + 5 = 0 by using a method different from the one you used in Part B. Show the steps of your work. (3 points)

OpenStudy (anonymous):

First step do you know how to find the discriminant

OpenStudy (anonymous):

No i dont.. :/

OpenStudy (anonymous):

Okay it is \[b ^{2}-4ac\]

OpenStudy (anonymous):

So you plug the numbers from your equation into this and solve... It would look like this: \[-16^{2}-4(9)(60)\]

OpenStudy (anonymous):

Okay so if i solve that that will give me the discriminant?

OpenStudy (anonymous):

I just evaluate that as much as a i can?

OpenStudy (anonymous):

@ihateschool18

OpenStudy (anonymous):

Yes it will tell me what you get so I can make sure you solve correctly

OpenStudy (anonymous):

Okay thank you so much so far :)

OpenStudy (anonymous):

okay so -2416

OpenStudy (anonymous):

Welcome :)

OpenStudy (anonymous):

okay so i solved the equation you gave me... It comes up to -2416

OpenStudy (anonymous):

One sec

OpenStudy (anonymous):

256-4(9)(60)= 256-2160= -1904

OpenStudy (anonymous):

So do you get part A?

OpenStudy (anonymous):

I understand a little. I have no idea what a discriminant does but as long as a i can do the work its fine. Thanks for part A

OpenStudy (anonymous):

@ihateschool18

OpenStudy (anonymous):

Yupp.. So moving to part B

OpenStudy (anonymous):

ohkayy.. So. I think you have to factor maybe.

OpenStudy (anonymous):

Good! Do you know how to factor this?

OpenStudy (anonymous):

ehhhh :/ i Could fiqure it out online without using your time. Could you help with c please?

OpenStudy (anonymous):

@ihateschool18

OpenStudy (anonymous):

Oh... Nvm its the same sortof.

OpenStudy (anonymous):

Ill just give it to you (2x-1)(2x+5). Are you sure you know what to do after that? Id you do we can move to c...

OpenStudy (anonymous):

If*

OpenStudy (anonymous):

It states in part A that there are multiple solutions, Does this mean they are asking for more then just the discriminate?

OpenStudy (anonymous):

i could probably figure it out thank you though :)

OpenStudy (anonymous):

Pare A was asking you to solve for the discriminant. After you solved for the discriminant the number you get tells you how many solutions you have.

OpenStudy (anonymous):

-_- i hate math.... okay.

OpenStudy (anonymous):

So if i just factor the other part that would be it for this problem?

OpenStudy (anonymous):

For part C because we factored we have to solve using a different method..I would probably use the quadratic formula or something...Do you know the formula?

OpenStudy (anonymous):

I could find it. I have it qritten down somewhere. and i know its online.

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b ^{2}-4ac}}{ 2a}\]

OpenStudy (anonymous):

Im sorry im super confused already.

OpenStudy (anonymous):

Why would i wanna use the quadiatric formula for this question? @ihateschool18

OpenStudy (anonymous):

Because we factored for part B it says in part C "Solve 2x2 -12x + 5 = 0 by using a method different from the one you used in Part B" so you cant use factoring to solve again

OpenStudy (anonymous):

oh.. Yes but why would this be the best method?

OpenStudy (anonymous):

Im not saying it would be the best really..I just find this the easiest. You can use whatever method you feel comfortable with

OpenStudy (anonymous):

Okay thanks!

OpenStudy (anonymous):

Welcome!